### v1

# Physically-based Lighting Augmentation for Robotic Manipulation

Abstract
Despite advances in data augmentation, policies trained via imitation learning still struggle to generalize across environmental variations such as lighting changes. To address this, we propose the first framework that leverages physically-based inverse rendering for lighting augmentation on real-world human demonstrations. Specifically, inverse rendering decomposes the first frame in each demonstration into geometric (surface normal, depth) and material (albedo, roughness, metallic) properties, which are then used to render appearance changes under different lighting. To ensure consistent augmentation across each demonstration, we fine-tune Stable Video Diffusion on robot execution videos for temporal lighting propagation. We evaluate our framework by measuring the structural and temporal consistency of the augmented sequences, and by assessing its effectiveness in reducing the behavior cloning generalization gap (40.1%) on a 7-DoF robot across 6 lighting conditions using 720 real-world evaluations. We further showcase three downstream applications enabled by the proposed framework.
Keywords: Robotic Manipulation, Lighting Augmentation, Imitation Learning
Figure 1: Our framework enables relighting of real-world human demonstrations through inverse rendering, improving imitation learning performance under diverse lighting conditions.
1 Introduction
Imitation learning offers a promising approach to deploying various robotic manipulation tasks using expert human demonstrations [1, 2, 3]. However, as pointed out by previous research [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], a critical challenge remains: the trained policies often struggle to generalize across variations in lighting, object texture and visual distractors. Capturing variations in these individual factors in real-world datasets is resource-intensive, and this challenge is further compounded when considering their combinations, making it difficult to address with real-world data collection alone.
Image augmentation techniques are widely adopted to improve generalization across environmental variations without requiring additional data collection. For example, domain randomization [14, 15] typically alters environmental parameters like backgrounds in simulation or applies visual perturbations like color jitter. Generative models have also been used to augment object textures and add visual distractors [4, 16, 17]. However, explicit modeling and manipulation of lighting in real-world demonstrations remains underexplored, largely due to the difficulty of recovering geometric and material information required for simulating accurate light–material interactions.
In this paper, we propose the first framework that applies physically-based inverse rendering for geometry- and material-aware lighting augmentation for imitation learning from real-world demonstrations in robotic manipulation. Our contributions are fourfold:
  • Modular Integration. We adopt an inverse rendering module [18] to decompose a single demonstration frame into geometric and material properties, which are then used by a rendering module to simulate new lighting on the decomposed properties.
  • Domain-Adapted Stable Video Diffusion. To ensure consistent augmentation across consecutive frames of each demonstration, we fine-tune Stable Video Diffusion (SVD) [19] for temporal lighting propagation. Fine-tuning data includes synthetic robot execution videos from Factor World [5] with varied lighting across tasks, and real-world execution videos from RoboNet [20] with visual degradation for radiometric accuracy.
  • Real-world Evaluation. We validate our framework by assessing structural and temporal consistency in augmented demonstrations, and through real-world experiments on a 7-DoF robot with an embodiment unseen during SVD fine-tuning. Under six varied lighting conditions, our method reduces the behavior cloning generalization gap by 40.1% across 720 evaluations on two tasks, compared to models trained without lighting augmentation.
  • Downstream Applications. We showcase augmentations on three additional environmental factors using geometric and material properties estimated by our framework.
2 Related Work
Data Augmentation for Robotic Manipulation.
Domain randomization [21, 22, 23, 24] has been widely studied to enhance policy robustness when transferring from simulation to the real world. Building upon this, recent efforts [4, 17, 16, 25, 26] have focused on semantic augmentation of real-world images leveraging text-driven generative models [27, 28] to introduce texture variation, visual distractors, etc. For example, ROSIE [17] proposes changing object textures by first segmenting augmentation regions [29] and then performing text-guided image inpainting [30]. Rendering techniques [31, 32, 33] have also been applied for viewpoint augmentation [34, 35, 36, 37]. Another line of work breaks long-horizon tasks into object-centric subtasks or manipulation skills and replays transformed demonstrations in simulation to generate new data from a limited number of examples [38, 39, 40, 41, 42]. In this paper, we focus on the underexplored problem of lighting augmentation on real-world human demonstrations for robotic manipulation.
Inverse Rendering and Relighting.
Relighting a scene typically requires identifying and altering its properties to produce the intended lighting effect [43]. Inverse rendering facilitates this by providing separate or joint estimations of geometry [44, 45, 46], material [47, 48, 49], and lighting [50, 51, 52] in a scene. Based on input requirements, inverse rendering can be categorized into single-view [53, 54, 55] and multi-view [56, 57, 58] methods. While methods such as DPI [59] and FIPT [60] produce high-fidelity relighting results, they rely on multi-view inputs for scene reconstruction and domain-specific datasets [61], making them incompatible with most existing robotic dataset camera setups. In this work, we present the first integration of single-view inverse rendering into robotic manipulation for explicit modeling of scene geometry and material properties.
3 Approach
3.1 Formulation and Overview
Given one episode of real-world human demonstration ($\mathbf{I}$, $\mathbf{P}$, $\mathbf{A}$) recorded from a fixed viewpoint, where $\mathbf{I}\in\mathbb{R}^{H\times W\times 3\times T}$ represents a sequence of $T$ RGB images, $\mathbf{P}\in\mathbb{R}^{p\times T}$ denotes proprioception states, and $\mathbf{A}\in\mathbb{R}^{a\times T}$ corresponds to actions, our goal is to perform physically-based lighting augmentation on $\mathbf{I}$ to produce a relit sequence $\mathbf{I}^{*}$ from the same viewpoint, while preserving structural and temporal consistency. Here, physically-based refers to physically-based rendering (PBR) for accurate light-material interactions.
Overview.
As shown in Fig. 2 the proposed framework consists of three components: (1) we begin by selecting the first frame $I_{0}$ from $\mathbf{I}$ and applying inverse rendering (Sec. 3.2) to estimate the geometric and material properties of the scene depicted in $I_{0}$; (2) the rendering module (Sec. 3.3) then uses the estimated properties and an environment map $E$ to relight $I_{0}$, producing $I_{0}^{*}$; and (3) the temporal propagation module (Sec. 3.4) propagates the lighting from $I_{0}^{*}$ across the full image sequence $\mathbf{I}$, resulting in the final augmented sequence $\mathbf{I}^{*}$. This forms an augmented episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$), which is then used to train the behavior cloning policy. The environment map $E$ has two use cases: (1) approximating the current lighting conditions to train a policy adapted to the environment (Sec. 4.2), and (2) introducing diverse lighting to contribute to a lighting-invariant policy (Sec. 5).
Figure 2: Given the first frame $I_{0}$ in a real-world human demonstration, the inverse rendering module estimates material ($A^{*}$, $R^{*}$, $M^{*}$) and geometric ($N^{*}$, $D^{*}$) properties. The rendering module uses these estimates and an environment map $E$ providing new lighting to produce a relit frame $I_{0}^{*}$. The new lighting is then propagated across the entire sequence $\mathbf{I}$, producing $\mathbf{I}^{*}$ and forming an augmented episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$) for behavior cloning.
3.2 Single-Frame Inverse Rendering
Following [18], we use a pre-trained network $\mathcal{P}$ to predict geometric and material properties of the scene depicted in $I_{0}$:
$A_{p},R_{p},M_{p},N_{p},D_{p}=\mathcal{P}(I_{0}),$ (1)
where $A_{p}$, $R_{p}$, and $M_{p}$ are predicted material properties (albedo, roughness, metallic), and $N_{p}$ and $D_{p}$ are predicted geometric properties (surface normal and depth). Examples of predicted properties can be found in Fig. 2. All predictions share the spatial resolution of $I_{0}$. Final property estimates are derived by minimizing the following objective:
$A^{*},R^{*},M^{*},N^{*},D^{*}=\arg\min_{A_{p},R_{p},M_{p},N_{p},D_{p}}\mathcal{L}_{p}(I_{0},I_{p}),$ (2)
where $I_{p}$ denotes the frame rendered using the predicted properties. $\mathcal{L}_{p}=\mathcal{L}_{\textit{re}}+\delta\mathcal{L}_{\textit{cons}}$, where $\mathcal{L}_{\textit{re}}$ denotes the reconstruction loss between the rendered frame $I_{p}$ and the original frame $I_{0}$, and $\mathcal{L}_{\textit{cons}}$ enforces consistency between the optimized and originally predicted properties, scaled by a factor $\delta$. Further details are provided in the supplementary material.
3.3 Single-Frame Relighting
To perform physically-based relighting, we render new lighting based on incoming radiance directions and modeled scene properties.
Light Sampling.
Given an environment map $E\in\mathbb{R}^{3}$, the incoming radiance $\lambda$ in direction $\omega_{i}$ is sampled as:
$\lambda(\omega_{i})=E(\omega_{i}),$ (3)
where $E$ is obtained either by approximating [18] or measuring [62, 63] the current lighting condition using graphics techniques, or by selecting from open-source HDRI datasets [64, 65].
Scene Modeling.
The estimated properties from Sec. 3.2 are used to model the material appearance $\mathcal{M}$ along direction $\omega_{i}$ of the scene depicted in $I_{0}$, using the widely adopted Disney BRDF [66]:
$\mathcal{M}(\omega_{i})=(1-M^{*})f_{\textit{diffuse}}(A^{*},R^{*},N^{*})+f_{\textit{specular}}(R^{*},N^{*}),$ (4)
where $f_{\textit{diffuse}}$ and $f_{\textit{specular}}$ model the diffuse and specular components of reflection, respectively; further details are provided in the supplementary material.
Scene Relighting.
Using the simplified rendering equation [67], the augmented frame $I_{0}^{*}$ under new lighting from environment map $E$ is computed as:
$I_{0}^{*}=\int_{\Omega}\mathcal{M}(\omega_{i})\lambda(\omega_{i})(\omega_{i}\cdot N^{*})\,\mathrm{d}\omega_{i},$ (5)
where $\Omega$ is the hemisphere centered at surface normal $N^{*}$, containing all incoming directions $\omega_{i}$. Note that once the property estimation for $I_{0}$ from Sec. 3.2 is completed, multiple augmentations can be applied using the same set of estimated properties.
3.4 Temporal Propagation
Domain-Adapted Stable Video Diffusion.
To meet the temporal consistency requirements of robotic manipulation, we adopt Stable Video Diffusion (SVD) [19] for its temporal modeling capabilities to propagate lighting from $I_{0}^{*}$ across the entire demonstration. We formulate this as a video-to-video translation task: lighting from the original image sequence $\mathbf{I}$ is transferred to generate the augmented sequence $\mathbf{I}^{*}$, guided by the reference frame $I_{0}^{*}$. This is defined as:
$\mathbf{I}^{*}=\delta(I_{0}^{*},\mathbf{I}_{\textit{input}}),\quad\mathbf{I}_{\textit{input}}=\text{Concatenate}(I_{0}^{*},\mathbf{I}_{1:T-1}),$ (6)
where $\delta$ denotes the SVD model.
Robotic Relighting Data Curation.
It’s important to note that the vanilla SVD underperforms in this task due to the absence of robotic elements like arms and grippers in its training data, resulting in a severely cartoonish appearance in generated robotic scenes. Following SVD’s original fine-tuning paradigm, we construct two datasets for different stages of fine-tuning:
Figure 3: (a) Examples of synthetic robot execution videos with lighting variations. (b) Examples of real-world robot execution videos with visual degradation.
  • Synthetic videos with lighting variation ($\mathcal{D}_{1}$, Fig. 3(a)). We generate domain-specific relighting data for robotic manipulation using the Factor World [5] benchmark, introducing lighting variation under identical task execution trajectories. All 42 built-in scenes covering 19 tasks (e.g., door opening) are used, each captured with six camera views: one fixed top-down and five randomized within constrained azimuth, inclination, and radius in the range $[-\pi/2,\pi/2]$ to ensure robot visibility. Each scene includes 30 lighting conditions by sampling ambient and diffuse RGB values from [25,255][25,255][ 25 , 255 ]. Videos are rendered at $\times 512$ 512 resolution with 100 frames at 24 fps. For fine-tuning, we randomly select pairs of videos showing the same task trajectories under different lighting, using one as the input $\mathbf{I}$ and the other as the target sequence $\mathbf{I}^{*}$.
  • Real-world videos with visual degradation ($\mathcal{D}_{2}$, Fig. 3(b)). Since existing real-world robotic datasets lack perfectly paired original and relit videos under identical task execution trajectories, we manually create video pairs by applying visual degradation on original videos from RoboNet [20] with 15 million video frames from 7 different robot platforms. We apply random visual transformations by sampling brightness, contrast, and saturation scaling factors from [0.2,1.9][0.2,1.9][ 0.2 , 1.9 ], and hue shifts from [0.5,0.5][-0.5,0.5][ - 0.5 , 0.5 ]. The degraded robot execution video serves as the input $\mathbf{I}$, while the original video acts as the output $\mathbf{I}^{*}$.
Fine-tuning.
We fine-tune SVD in two stages. In Stage 1, we use synthetic videos from $\mathcal{D}_{1}$, which provide the main data for domain adaptation and reduce overfitting via controlled lighting variation. However, due to the non-physically-based lighting in Factor World, this stage lacks radiometric accuracy, motivating fine-tuning on real-world videos from $\mathcal{D}_{2}$ in Stage 2. Fine-tuning is performed on a pretrained 14-frame SVD model using a single H100 GPU for 8,000 steps for Stage 1 and 1,000 steps for Stage 2. A limited number of steps are applied to Stage 2 due to the absence of ground-truth relit pairs. Further details and ablation studies are provided in the supplementary material.
4 Experiments
4.1 Visual Lighting Quality Evaluation
We apply lighting augmentation to 10 randomly selected sequences (600 frames) from the real-world human demonstrations in Sec.4.2 using our method and IC-Light [68], a widely recognized text-prompt-based relighting method. Ground-truth episodes are recorded under a side-mounted blue LED panel light under the same task execution trajectories. Two sets of metrics are adopted: (1) LPIPS [69] and SSIM [70] are computed between the ground-truth and relit image sequences to assess structural consistency and fidelity. (2) Temporal LPIPS and Temporal SSIM are computed between consecutive frames in the relit sequences to assess temporal consistency. Since IC-Light does not support video relighting, we apply it frame-by-frame using a fixed random seed and the prompt “blue LED light” for consistency. We perform relighting using an environment map approximated from the ground truth episodes with [18]. Qualitative and quantitative evaluation results can be found in Fig. 4. Our method achieves better performance by estimating scene-level properties, which are more compatible with robotic settings, whereas IC-Light’s foreground/background separation is better suited for human relighting. Our temporal propagation module further improves visual consistency by reducing abrupt changes between frames.
Method LPIPS ($\downarrow$) SSIM ($\uparrow$) Temporal LPIPS ($\downarrow$) Temporal SSIM ($\uparrow$)
IC-Light 0.5274 0.5359 0.121 0.898
Ours 0.3269 0.7351 0.035 0.978
Figure 4: Qualitative (top) and quantitative (bottom) evaluation of lighting augmentation. Although exceeding the baseline, the lighting in our augmented episodes appears more diffused compared to the ground truth, as indicated by weaker side shadows. This discrepancy mainly results from the rough environment map approximation (32×1632$\times$ 1632 × 16 resolution) without specialized graphics equipment.
4.2 Real-world Evaluation
Lighting Setup.
White light with minimal shadows (Ambient, Fig. 5(a)) is used during expert human demonstration collection for policy training. During evaluation, six lighting conditions are designed using a programmable RGB LED panel. This includes two sets with different evaluation purposes: (1) Daily-use conditions — Side (strong shadows), Regional (blue regional highlight), and Dim (low-light); (2) Artificial RGB lighting — Red, Green, and Blue — with illuminance levels of 4,600 lux and RGB values (255, 0, 0), (0, 255, 0), and (0, 0, 255), respectively, designed to evaluate each policy under single-source lighting. These colors are chosen based on the principle of light transport linearity, which allows any colored illumination to be reconstructed as their linear combination. Examples are shown in Fig. 5(b).
Manipulation Tasks.
Pear: Grasp a randomly placed green plastic pear using a vacuum gripper and place it on a plate. Square: Grasp a randomly placed black square nut by its handle using a parallel jaw gripper and insert it into a matching red square peg. To better assess performance under challenging lighting conditions, we report performance separately for the reaching (R) subtask, which involves reaching the area of the pear or square nut, and the pick-and-place (PnP) outcome, defined as successfully picking up the object and placing it in the corresponding plate or peg.
Figure 5: (a) Experimental setup under Ambient lighting. (b) The six lighting conditions used for generalization gap evaluation. (c) Objects used in the two manipulation tasks.
Implementation Details.
We perform behavior cloning using ResNet18 [71] and Multi-Layer Perceptron (BC-MLP) to map features to actions. Random cropping (Crop) is used as the default augmentation due to its widespread use and proven effectiveness [5, 39]. We evaluate two variants: +Jitter, which adds color jitter, and +Ours, which adds our proposed lighting augmentation. Color jitter is selected as a baseline due to its ability to mitigate the impact of color and illumination variation. Training and evaluation are conducted on a 7-DoF UFactory XArm7 robot observed by a single externally mounted RGB camera. Each task uses 200 human demonstrations to train the baseline policies Crop and +Jitter. For +Ours, 10 demonstrations are randomly selected and augmented with six environment maps approximating the six evaluation lighting conditions, producing 260 training episodes per task for policy training. See the supplementary material for full implementation details.
Evaluation Protocol.
For each combination of task $\tau$, lighting condition L, and augmentation method Aug, we conduct 20 real-world evaluations. The success rate for each combination is denoted as $S_{\textit{L}}^{\textit{Aug}}(\tau)$. We use BC-MLP trained and evaluated under Ambient lighting with Crop augmentation as the base policy. The corresponding subtask success rates are: $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Pear}\text{-}\textit{R})$ = 1.00, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Pear}\text{-}\textit{PnP})$ = 0.60, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Square}\text{-}\textit{R})$ = 1.00, $S_{\textit{Ambient}}^{\textit{Crop}}(\textit{Square}\text{-}\textit{PnP})$ = 0.85. Performance degradation under unseen evaluation lighting conditions is measured by the Generalization Gap, defined as the normalized drop in success rate relative to the base policy:
$\textit{Generalization Gap}=(S_{\textit{Ambient}}^{\textit{Crop}}(\tau)-S_{\textit{L}}^{\textit{Aug}}(\tau))/S_{\textit{Ambient}}^{\textit{Crop}}(\tau).$ (7)
Results.
Tab. 1 presents the generalization gap of BC-MLP under six unseen lighting conditions across 720 real-world evaluations. Averaged over both subtasks across the two tasks, our method outperforms Crop by 40.1% and +Jitter by 35.7%. In the daily-use lighting set, Crop shows a major drop in performance, except under Dim. While +Jitter performs reasonably on the reaching subtask, it consistently fails in the final pick-and-place outcome. +Ours improves performance across both subtasks under these conditions. In the artificial RGB lighting set, both Crop and +Jitter fail in most cases, even on the reaching subtask. Interestingly, +Jitter displays strong preferences for different RGB lighting in the two tasks, likely due to its use of global color shifts. Our method maintains strong reaching performance but struggles with pick-and-place in these more extreme lighting conditions. Common failure modes for Crop and +Jitter involve moving directly to the plate or peg without interacting with the pear or nut. Failures in +Ours often involve reaching the pear or nut with the gripper open and hovering just above the object, without executing a pick action. This may relate to the altered object appearance under high-intensity lighting (4,600 lux), where high reflectivity could disrupt perception. See the limitations section for further details and the supplementary video for examples of common failure modes across different methods.
Table 1: Generalization Gap ($\downarrow$) with different augmentation methods under unseen lighting. The best-performing policy is in bold, and the second-best is underscored.
Lighting Crop +Jitter +Ours
Pear Square Pear Square Pear Square
R PnP R PnP R PnP R PnP R PnP R PnP
Side 1.00 1.00 1.00 1.00 0.70 1.00 0.85 1.00 0.92 1.00 0.10 0.53
Regional 0.95 1.00 1.00 1.00 0.55 1.00 0.50 1.00 0 0.67 0 0.47
Dim 0 0.16 0.05 0.35 0.1 0.83 0.05 0.47 0 0.34 0 0.06
Red 1.00 1.00 1.00 1.00 0.70 1.00 0.90 1.00 0.40 1.00 0.10 1.00
Green 0.85 1.00 0.90 1.00 1.00 1.00 0.95 1.00 0 1.00 0 0.88
Blue 1.00 1.00 0.95 1.00 1.00 1.00 0.55 1.00 0.15 0.92 0.05 1.00
4.3 Ablation Study
Figure 6: Ablation study: (a) effect of the number of lighting conditions; (b) effect of the number of augmented episodes.
Impact of the Number of Lighting Conditions for Augmentation.
We perform augmentation using 1 (Red), 3 (Red, Green, Blue), 6 (all conditions from Sec. 4.2), and 9 lighting conditions (6 plus three additional lights with RGB values (255, 255, 0), (255, 0, 255), and (0, 255, 255)). Evaluation is performed in single-source Red light to control for lighting composition effects and isolate augmentation impact. Reaching success rates on Square task, averaged over 10 trials, are shown in Fig. 6(a). Including the currently observed lighting (Red) in augmentation immediately boosts performance (from 0%). However, performance degrades when too many unrelated lighting conditions are added, likely due to the reduced presence of target augmentations (Red) in the expanded set. See Sec. 7 for further discussion.
Impact of the Number of Episodes for Augmentation.
We select 10, 25, 75, and 100 episodes from 200 demonstrations and augment each with six lighting conditions. Evaluation follows the same setup as above. As shown in Fig. 6(b), performance is relatively unaffected by the number of augmented episodes, with no signs of overfitting — possibly due to the lighting diversity within each episode. To save computation, we use the minimal setting of 10 episodes for augmentation.
5 Applications
Through scene decomposition via inverse rendering, we extend our framework to augment three additional environmental factors (background, object texture, and distractors) beyond lighting (Fig. 7(a)) on existing datasets, as demonstrated on BridgeData v2 [72].
5.1 Background Augmentation
Background appearance is tied to global illumination; therefore, we render segmented scene geometry with different environment maps to create new backgrounds with corresponding lighting. Scene geometry is reconstructed by triangulating the depth map $D^{*}$ from Sec. 3.2 into a mesh $\Lambda$:
$\Lambda=\text{Triangulate}\left(K^{-1}[x\;\;y\;\;1]^{\top}D^{*}(x,y)\right),$ (8)
[…]
Figure 7: Examples of applications: (a) Lighting augmentation with different environment maps; (b) Background augmentation with different environment maps; (c) Texture augmentation via albedo adjustment; (d) Gravity-based distractor placement, with distractor mesh sourced from [74].
5.2 Object Texture Augmentation
Robotic tasks like grasping and pushing rely on consistent object geometry and frictional properties  [75, 76]. We address this by applying object texture augmentation with preserved visual roughness and geometry. Following the mesh segmentation in Sec. 5.1, we segment the object mesh and render with altered material properties. Specifically, albedo $A^{*}$ sets the base color, roughness $R^{*}$ controls surface scattering, and metallic $M^{*}$ defines the metallic effect. By adjusting only the albedo $A^{*}$ while keeping roughness $R^{*}$ and metallic $M^{*}$ fixed, and reapplying Eq. 4, we achieve object texture augmentation with consistent visual roughness and geometry. Examples are shown in Fig. 7(c).
5.3 Distractor Placement in Cluttered Environments
Previous methods typically place distractors via inpainting using text-driven generative models [17, 16, 4]. We take a different approach by leveraging the meshes of the scene and distractors for physically plausible placement in Blender [77]. As shown in Fig. 7(d), we first align the surface normal with Blender’s default gravity axis (Z-axis). This requires the user to manually select a mesh face whose surface normal is parallel to the gravity direction using the Blender API. A random position within the bounding box of $\Lambda$ (Eq. 8) is sampled, and the distractor mesh — assigned a pseudo-mass of 1kg at its geometric center — is dropped to simulate its motion under gravity. The final placement of the distractor is determined by its stable resting state following the gravity-based simulation. The full simulation process is available in the supplementary video.
6 Conclusion
We propose the first framework that applies physically-based lighting augmentation on real-world human demonstrations for robotic manipulation. By decomposing robotic scenes into geometric and material properties, we perform lighting augmentation on a single frame and propagate it across the entire demonstration using finetuned Stable Video Diffusion. We validate our framework through qualitative and quantitative visual quality evaluations, 720 real-world trials under six varied lighting conditions using a 7-DoF robot, and complementary ablation studies. We further showcase augmentations on three additional environmental factors enabled by our framework. This work marks an initial step toward integrating physical realism into visual augmentation, with the potential to assist future research on improving imitation learning generalization under environmental variation.
7 Limitations and Future work
Environment Lighting Approximation.
Although we provide physically-based lighting augmentation by simulating light-material interactions, the absence of an environment map that perfectly reconstructs the currently observed lighting leads to reduced performance. One example is shown in Fig. 8, where under high-intensity red LED lighting (4600 lux), the plastic pear and plate exhibit strong reflections, differing from our augmentation. The discrepancy results from the side-mounted high-intensity LED producing strong directional lighting, while the approximated environment map failed to capture this and provides primarily diffuse illumination with a shifted color tone. This is partially due to the limited precision of the single-view-based environment map approximation (32×1632$\times$ 1632 × 16 resolution). Capturing current lighting using graphical equipment [62] to generate accurate environment maps could mitigate this issue. Future work will focus on improving lighting approximation and deriving high-precision environment maps from single-view input for robotic scenes.
Figure 8: Failure cases observed under high-intensity lighting.
Efficiency.
The current temporal propagation module only applies to lighting augmentation and does not extend to the other three environmental factors discussed in Sec. 5, as they are not captured by the curated SVD fine-tuning dataset. Consequently, augmentation for the additional factors is currently performed frame by frame, which makes the proposed applications more suitable for offline augmentation of human demonstration data. Future work will focus on improving efficiency through selective optimization of geometric and material properties, taking into account environmental similarities in robotic data collection settings.
A Universal Lighting Composition.
This paper focuses on using the proposed lighting augmentation framework to improve performance under currently observed lighting conditions, as the specific and quantified effects of different lighting on robot learning remain largely unknown. Future work will aim to systematically study how various lighting conditions impact policy performance and, based on this understanding, develop effective augmentation strategies for constructing a universal lighting composition to support lighting-invariant policies.
Visual-Physical Gap.
The term physically-based in this paper refers to physically-based rendering. While we follow physically accurate light–material interactions from a computer graphics perspective, these effects are purely visual and do not affect motion dynamics, which is a critical aspect of robotic manipulation.